A fossilized leaf contains 12% of its normal amount of carbon 14. How old is the fossil (to the nearest year)? Use 5600 years as the half-life of carbon 14.
Answer by fractalier(6550)

We use the equation
A(t) = Ao*e^(-t/5600)
Here, A(t) = .12Ao so
.12Ao = Ao*e^(-t/5600)
.12 = e^(-t/5600)
Take the ln of both sides and get
-2.12 = -t/5600
t = 11873 years